A rectangle is drawn so the width is 31 inches longer than the height. If the rectangle's diagonal measurement is 41 inches, find the height.
inches =

Answer by ankor@dixie-net.com(22740)

A rectangle is drawn so the width is 31 inches longer than the height.
If the rectangle's diagonal measurement is 41 inches, find the height.
Using pythag
h^2 + (h+31)^2 = 41^2
FOIL (h+31)(h+31)
h^2 + h^2 + 31h + 31h + 961 = 1681
combine like terms
2h^2 + 62h + 961 - 1681 = 0
2h^2 + 62h - 720 = 0
simplify, divide by 2
h^2 + 31h - 360 = 0
Factors to
(h + 40)(h - 9) = 0
The positive solution is all we want here
h = 9 inches is the height
:
:
Check; the other side: 9+31
9^2 + 40^2 = 41^2